Find the maclaurent series of arctan x
1
Expert's answer
2019-04-03T10:05:25-0400

To obtain the maclaurent series of arctan x, use the standard integral

"\arctan x=\int_{0}^{x}{\frac{dt}{1+{{t}^{2}}}}"

Consider the sum of a geometric series

"\frac{1}{1+{{t}^{2}}}=\frac{1}{1-\left( -{{t}^{2}} \right)}=1-{{t}^{2}}+{{t}^{4}}-{{t}^{6}}+..."

Substituting this sum into the integral, we obtain

"\int_{0}^{x}{\frac{dt}{1+{{t}^{2}}}}=\int_{0}^{x}{\left( 1-{{t}^{2}}+{{t}^{4}}-{{t}^{6}}+... \right)dt}"

"=\left( t-\frac{{{t}^{3}}}{3}+\frac{{{t}^{5}}}{5}-\frac{{{t}^{7}}}{7}+... \right)_{0}^{x}=x-\frac{{{x}^{3}}}{3}+\frac{{{x}^{5}}}{5}-\frac{{{x}^{7}}}{7}+..."

Thus

"\arctan x=x-\frac{{{x}^{3}}}{3}+\frac{{{x}^{5}}}{5}-\frac{{{x}^{7}}}{7}+..."

or


"\arctan x=\sum\limits_{n=0}^{\infty }{{{\left( -1 \right)}^{n}}\frac{{{x}^{2n+1}}}{2n+1}}"

for

"\left| x \right|\le 1"

This is the Maclaurent series of arctan x.